An infinitely long conductor PQR is bent to form a right angle. A current I flows at the point M is H1.Now,another infinitely long straight conductor QS is connected at Q so that current I/2 in QR as well as in QS, the current in PQ remaining unchanged. Magnetic field at M is now H2. Ratio of H1/H2 is... (a) 1/2 (b) 1 (c) 2/3 (d) 2 fig : M P--------Q S | | R


Grade:upto college level

1 Answers

Shubham Kumar Gupta Ggps
6 years ago
H1=ui/4πrH2=u(I/2)/4πr + ui/4πrH1/H2=2/3 As due to QR magnetic field is 0And simply apply the formula